# A congruence subgroup property for symmetric mapping class groups

### v1

@@ 4. The congruence subgroup property for relative mapping class groups
Let us observe that, for $\partial S\neq\emptyset$, the natural representation ${\Gamma}(S,\partial S)\to\operatorname{Out}(\Pi)$ is not faithful. Hence, the congruence subgroup property, as formulated in Section 1 does not make sense for relative mapping class groups. For this reason, in [7, Section 2.7], it was introduced the following construction.
Let ${\sigma}:=\{\delta_{1},\ldots,\delta_{k}\}$ be the set of connected components of $\partial S$ and let $\widetilde{S}$ be the surface (without boundary) obtained from $S$ attaching to each curve $\delta_{i}$ a $2$-punctured closed disc $D_{i}$, for $i=1,\dots k$. Let then ${\Gamma}(\widetilde{S})_{\sigma}$ be the stabilizer in ${\Gamma}(\widetilde{S})$ of the set of isotopy classes of the curves contained in ${\sigma}$. The relative mapping class group ${\Gamma}(S,\partial S)$ acts on the set ${\sigma}$ and the associated representation in the symmetric group ${\Gamma}(S,\partial S)\to\Sigma_{k}$ is surjective. There is then a natural isomorphism:
(5) ${\Gamma}(\widetilde{S})_{\sigma}\cong(\Sigma_{2})^{k}\rtimes{\Gamma}(S,\partial S),$
where the mapping class group ${\Gamma}(S,\partial S)$ acts on the product of symmetric groups $(\Sigma_{2})^{k}$ by permuting its factors through the epimorphism ${\Gamma}(S,\partial S)\to\Sigma_{k}$ defined above.
The isomorphism (5) identifies the relative mapping class group ${\Gamma}(S,\partial S)$ with a subgroup of the mapping class group ${\Gamma}(\widetilde{S})$. We then define the congruence topology on ${\Gamma}(S,\partial S)$ as the topology induced, via this embedding, by the congruence topology of ${\Gamma}(\widetilde{S})$. The relative procongruence mapping class group ${\check{\Gamma}}(S,\partial S)$ is the associated completion of ${\Gamma}(S,\partial S)$.
@@
For a finite group acting faithfully on $S$, we let ${\Gamma}_{G}(S,\partial S)$ be the centralizer of $G$ in ${\Gamma}(S,\partial S)$. We define the congruence topology on ${\Gamma}_{G}(S,\partial S)$ as the one induced by the congruence topology of ${\Gamma}(S,\partial S)$ and we denote by ${\check{\Gamma}}_{G}(S,\partial S)$ the associated completion. We can then formulate the corresponding congruence subgroup problem for ${\Gamma}_{G}(S,\partial S)$.
There is a natural short exact sequence:
@@
From the above two exact sequences, it follows, in particular, that the congruence subgroup property for the mapping class group ${\Gamma}(S)$ implies the same property for the relative mapping class group ${\Gamma}(S,\partial S)$. A similar argument yields the following generalization of Theorem A and Corollary B:
@@
Let $S$ be an oriented connected hyperbolic surface of finite type and $G$ a finite group acting faithfully on $S$. If the congruence subgroup property holds for the mapping class group ${\Gamma}(S_{/G}^{\circ})$ of the surface $S_{/G}^{\circ}$, then it holds for the centralizer ${\Gamma}_{G}(S,\partial S)$ of $G$ in the relative mapping class group ${\Gamma}(S,\partial S)$. In particular, for $g(S_{/G})\leq 2$, there is a natural isomorphism ${\widehat{\Gamma}}_{G}(S,\partial S)\cong{\check{\Gamma}}_{G}(S,\partial S)$.
@@
$1\to(\prod_{i=1}^{k}\tau_{{\delta}_{i}}^{\mathbb{Z}})^{G}\to{\Gamma}_{G}(S,\partial S)\to{\Gamma}_{G}(S)\to 1$
@@
$1\to(\prod_{i=1}^{k}\tau_{{\delta}_{i}}^{\widehat{\mathbb{Z}}})^{G}\to{\check{\Gamma}}_{G}(S,\partial S)\to{\check{\Gamma}}_{G}(S)\to 1,$
where the subgroup of invariants $(\prod_{i=1}^{k}\tau_{{\delta}_{i}}^{\widehat{\mathbb{Z}}})^{G}$ identifies with the closure of the subgroup $(\prod_{i=1}^{k}\tau_{{\delta}_{i}}^{\mathbb{Z}})^{G}$ in ${\check{\Gamma}}_{G}(S,\partial S)$. Theorem 4.1 then follows from the two above exact sequences and Theorem A if we show that the profinite group $(\prod_{i=1}^{k}\tau_{{\delta}_{i}}^{\widehat{\mathbb{Z}}})^{G}$ is the profinite completion of the group $(\prod_{i=1}^{k}\tau_{{\delta}_{i}}^{\mathbb{Z}})^{G}$. But this is clear since the profinite completion of a finitely generated abelian group induces the profinite completion on each of its subgroups. ∎
5. Proof of Theorem C
@@
    the closure of the normalizer $N_{{\Gamma}(S,\partial S)}(G)$ in ${\widehat{\Gamma}}(S,\partial S)$ coincides with the profinite completion of $N_{{\Gamma}(S,\partial S)}(G)$.
Since ${\Gamma}(S)$ is a good group for $g(S)\leq 2$, from Lyndon-Hochschild-Serre spectral sequence and the short exact sequences (6) and (7), we deduce that the relative mapping class group ${\Gamma}(S,\partial S)$ is also good for $g(S)\leq 2$.
There is a version of the short exact sequence (2) for relative mapping class groups:
$1\to G\to N_{{\Gamma}(S,\partial S)}(G)\to{\Gamma}(S_{/G}^{\circ},\partial S_{/G}^{\circ})_{K}\to 1$
By the previous remark, this implies that $N_{{\Gamma}(S,\partial S)}(G)$ is a good group for $g(S_{/G})\leq 2$. Theorem 4.1 then implies in particular that, for $g(S_{/G})\leq 2$, the subspace topology on $N_{{\Gamma}(S,\partial S)}(G)\subseteq{\Gamma}(S,\partial S)$, induced by the profinite topology of ${\Gamma}(S,\partial S)$, is the full profinite topology on this group.

### v2

@@ 4. The congruence subgroup property for relative mapping class groups
Let us observe that, for $\partial S\neq\emptyset$, the natural representation ${\Gamma}(S,\partial S)\to\operatorname{Out}(\Pi)$ is not faithful. Hence, the congruence subgroup property, as formulated in Section 1 does not make sense for relative mapping class groups. For this reason, in [7, Section 2.7], we introduced the following construction.
[…]
The relative mapping class group ${\Gamma}(S,\partial S)$ and the product of relative mapping class groups $\prod_{i=1}^{k}{\Gamma}(D_{i},\partial D_{i})$ then identify, respectively, with a subgroup and a normal subgroup of the stabilizer ${\Gamma}(\widetilde{S})_{\sigma}$ and there holds:
(5) ${\Gamma}(\widetilde{S})_{\sigma}=(\prod_{i=1}^{k}{\Gamma}(D_{i},\partial D_{i}))\cdot{\Gamma}(S,\partial S).$
[…]
@@
For a finite group acting faithfully on $S$, we let ${\Gamma}(S,\partial S)^{G}$ be the centralizer of $G$ in ${\Gamma}(S,\partial S)$. We define the congruence topology on ${\Gamma}(S,\partial S)^{G}$ as the one induced by the congruence topology of ${\Gamma}(S,\partial S)$ and we denote by ${\check{\Gamma}}(S,\partial S)^{G}$ the associated completion. We can then formulate the corresponding congruence subgroup problem for ${\Gamma}(S,\partial S)^{G}$.
Let us observe that there is a natural short exact sequence:
@@
We then have the following generalization of Theorem A and Corollary B:
@@
Let $S$ be an oriented connected hyperbolic surface of finite type and $G$ a finite group acting faithfully on $S$. If the congruence subgroup property holds for the mapping class group ${\Gamma}(S_{/G}^{\circ})$ of the surface $S_{/G}^{\circ}$, then it also holds for the centralizer ${\Gamma}(S,\partial S)^{G}$ of $G$ in the relative mapping class group ${\Gamma}(S,\partial S)$. In particular, for $g(S_{/G})\leq 2$, there is a natural isomorphism ${\widehat{\Gamma}}(S,\partial S)^{G}\cong{\check{\Gamma}}(S,\partial S)^{G}$.
@@
$1\to(\prod_{i=1}^{k}\tau_{{\delta}_{i}}^{\mathbb{Z}})^{G}\to{\Gamma}(S,\partial S)^{G}\to{\Gamma}(S)^{G}\to 1.$
@@
$1\to(\prod_{i=1}^{k}\tau_{{\delta}_{i}}^{\widehat{\mathbb{Z}}})^{G}\to{\check{\Gamma}}(S,\partial S)^{G}\to{\check{\Gamma}}(S)^{G}\to 1,$
where the subgroup of invariants $(\prod_{i=1}^{k}\tau_{{\delta}_{i}}^{\widehat{\mathbb{Z}}})^{G}$ identifies with the closure of the subgroup $(\prod_{i=1}^{k}\tau_{{\delta}_{i}}^{\mathbb{Z}})^{G}$ in ${\check{\Gamma}}(S,\partial S)^{G}$. Theorem 4.1 then follows from the two above exact sequences and Theorem A if we show that the profinite group $(\prod_{i=1}^{k}\tau_{{\delta}_{i}}^{\widehat{\mathbb{Z}}})^{G}$ is the profinite completion of the group $(\prod_{i=1}^{k}\tau_{{\delta}_{i}}^{\mathbb{Z}})^{G}$. But this is clear since the profinite completion of a finitely generated abelian group induces the profinite completion on each of its subgroups. ∎
5. Conjugacy separability
We will prove the following more general version of Theorem C:
Theorem 5.1.
For $S$ an oriented connected hyperbolic surface of finite type of genus $\leq 2$, we have:
  1. (i)
    The natural monomorphism ${\Gamma}(S,\partial S)\hookrightarrow{\widehat{\Gamma}}(S,\partial S)$ induces a bijection between conjugacy classes of solvable finite subgroups. In particular, solvable finite subgroups of ${\Gamma}(S,\partial S)$ are subgroup conjugacy distinguished.
  2. (ii)
    Torsion elements of ${\Gamma}(S,\partial S)$ are conjugacy distinguished.
  3. (iii)
Proof.
@@ 5. Conjugacy separability
    the closure of the normalizer $N_{{\Gamma}(S,\partial S)}(G)$ in ${\widehat{\Gamma}}(S,\partial S)$ identifies with the profinite completion of $N_{{\Gamma}(S,\partial S)}(G)$.
The claim that ${\Gamma}(S,\partial S)$ is good for $g(S)\leq 2$ follows from the fact that ${\Gamma}(S)$ is a good group for $g(S)\leq 2$ and from the Lyndon-Hochschild-Serre spectral sequence applied to the natural homomorphism from the short exact sequence (6) to the short exact sequence (7).
Let us then prove that the normalizer $N_{{\Gamma}(S,\partial S)}(G)$ is also a good group for $g(S_{/G})\leq 2$ (and so, in particular, for $g(S)\leq 2$). Let us observe first that there is the following version of the short exact sequence (2) for relative mapping class groups:
(8) $1\to G\to N_{{\Gamma}(S,\partial S)}(G)\to{\Gamma}(S_{/G}^{\circ},\partial S_{/G}^{\circ})_{K}\to 1.$
Since ${\Gamma}(S_{/G}^{\circ},\partial S_{/G}^{\circ})_{K}$ is a finite index subgroup of ${\Gamma}(S_{/G}^{\circ},\partial S_{/G}^{\circ})$ and the latter, as we proved above, is good for $g(S_{/G})\leq 2$, from Shapiro’s lemma, it follows that ${\Gamma}(S_{/G}^{\circ},\partial S_{/G}^{\circ})_{K}$ is also good for $g(S_{/G})\leq 2$. By the Lyndon-Hochschild-Serre spectral sequence applied to the natural homomorphism which maps the short exact sequence (8) to its profinite completion, we then have that $N_{{\Gamma}(S,\partial S)}(G)$ is a good group for $g(S_{/G})\leq 2$.
By Theorem 4.1, for $g(S_{/G})\leq 2$, the closure of the centralizer ${\Gamma}(S,\partial S)^{G}$ in ${\widehat{\Gamma}}(S,\partial S)$ identifies with the profinite completion of ${\Gamma}(S,\partial S)^{G}$. Since ${\Gamma}(S,\partial S)^{G}$ is a finite index subgroup of the normalizer $N_{{\Gamma}(S,\partial S)}(G)$, it follows that, for $g(S_{/G})\leq 2$ (and so, in particular, for $g(S)\leq 2$), the closure of the normalizer $N_{{\Gamma}(S,\partial S)}(G)$ in ${\widehat{\Gamma}}(S,\partial S)$ identifies with the profinite completion of $N_{{\Gamma}(S,\partial S)}(G)$. ∎